OpenStudy (anonymous):

If the following system of equations was written as a matrix equation in the form AX = C, and matrix A was expressed in the form [a c b d] , find the value of a - b + c + d.

OpenStudy (anonymous):

2x+8y=7 4x-2y=9

OpenStudy (anonymous):

matrix A is expressed in the form \[\left[\begin{matrix}a & b \\ c & d\end{matrix}\right]\] is that correct?

OpenStudy (anonymous):

I need to know if it's like i said or if it's in the form of [a c b d]

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Its the first one a c b d

OpenStudy (anonymous):

AX = C \[\left[\begin{matrix}2 & 8 \\ 4 & 2\end{matrix}\right]\left(\begin{matrix}x \\ y\end{matrix}\right)=\left(\begin{matrix}7 \\ 9\end{matrix}\right)\]

OpenStudy (anonymous):

can u see that a=2, c=8, b=4 and d=2? now u can calculate a-b+c+d, try it yourself.

OpenStudy (anonymous):

its not -2 on the bottom right ?

OpenStudy (anonymous):

oh yeh it is, sorry i'm blind.

OpenStudy (anonymous):

No its alright

OpenStudy (anonymous):

So it would be 4 right

OpenStudy (anonymous):

yep.

OpenStudy (anonymous):

Thankss

OpenStudy (anonymous):

no problem brook! I think the key to solving these type of questions is knowing how to put the equations into a matrix, u do know how to do it now right?

OpenStudy (anonymous):

Nopp but I just found out, thank youu

OpenStudy (anonymous):

ok, that's good. no problem.